one of the two digits of a two digit number is thrice the other number.If the two digits are interchanged and the resultant is added to the original number we get 88.find the original number?

Answer by lwsshak3(11628)

one of the two digits of a two digit number is thrice the other number.If the two digits are interchanged and the resultant is added to the original number we get 88.find the original number?
***
let u=units digit
let t=tens digit
u=3t
..
10u+t+10t+u=88
30t+t+10t+3t=88
44t=88
t=2
u=3t=6
original number=10t+u=20+6=26